what volume would 120.0g of nitrogen gas occupy at 59.50 c and 0.566 atm


1
Expert's answer
2022-01-10T13:56:52-0500

59.50oC = 59.50 + 273.15 = 332.65 K


"n(N_2)=\frac{120.0\ g}{28.01\ g/mol}=4.284\ mol"


"R=0.08206\frac{L\cdot{atm}}{mol\cdot{K}}"


According to the ideal gas law,

"PV=nRT" ;

"V=\frac{nRT}{P}=\frac{4.284\ mol\times0.08206\frac{L\cdot{atm}}{mol\cdot{K}}\times332.65\ K}{0.566\ atm}=207\ L"


Answer: 207 L